OpenStudy (anonymous):

The area of a rectangle with a length of 8 inches is 48 square inches. What is the perimeter of the rectangle?

OpenStudy (anonymous):

Area = length x width. 48 = 8w. w = 6. Perimeter = 2l + 2w = 2(8) + 2(6) = 16 + 12 = 28.

OpenStudy (anonymous):

Area = length x width. 48= 8w

OpenStudy (anonymous):

fma